OpenStudy (anonymous):

Use the three steps to solve the problem. One number is 800 less than 3 times the other, and their sum is 1,200. Find the numbers.

OpenStudy (anonymous):

One number is x. The other number is expressed in terms of this number, in terms of x. So, the other number is 3x - 800. Add both together and set to 1200. Solve for x.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Where are you stuck?

OpenStudy (anonymous):

i dont understand none of it :(

OpenStudy (anonymous):

Ok, from the very start. You are being asked to add two numbers and get 1200. So, an initial equation could be x + y = 1200. This follows follows from "...and their sum is...". Do you follow?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

One number is x and the other is y, but you can express y in terms of x.

OpenStudy (anonymous):

let x be the number then less than means subtract i think 800-3x=0......1 800+3x=1200...........2 solve further

OpenStudy (anonymous):

As I pointed out in the first post, the second number, y, is 3x - 800.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So, add x and 3x - 800 and set that = 1200.

OpenStudy (anonymous):

262

OpenStudy (anonymous):

938

OpenStudy (anonymous):

Well, you have x + (3x - 800) = 1200 -> (x + 3x) - 800 = 1200 -> 4x - 800 = 1200. Can you find x from here?

OpenStudy (anonymous):

Add 800 to each side.

OpenStudy (anonymous):

idk this is to confusing

OpenStudy (anonymous):

Are you having trouble adding 800 to each side of the equation 4x - 800 = 1200?

OpenStudy (anonymous):

ik how to add just this problem is so confusing

OpenStudy (anonymous):

200??

OpenStudy (anonymous):

Any problem can be hard if there are a lot of steps and you desire to go from step 1 right to step 10 or step 15. Problems definitely do get easier when you go only to the next step. You will eventually get there, but it isn't always quick. In the beginning, you have to do all the steps, but then it gets easier because you can then either skip a couple steps or just do them faster with practice.

OpenStudy (anonymous):

4x - 800 = 1200 -> 4x = 2000 -> x = 500. So 500 is one of your numbers. Can you get the other number?

OpenStudy (anonymous):

Hint, they add to 1200. So, 500 + y = 1200.

OpenStudy (anonymous):

okay 700

OpenStudy (anonymous):

Yes, you got it! It is helpful to go back to the original problem and then you will be able to see how those numbers satisfy both conditions.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Before you go, one question.

OpenStudy (anonymous):

ya?

OpenStudy (anonymous):

Do you find that you are getting this even a little better now? The whole goal is for students to learn a little bit and feel more comfortable with math. Too often some go throughout life with an uneasiness about math, and they wish they didn't.

OpenStudy (anonymous):

yes i think i am getting a little bit better

OpenStudy (anonymous):

The solving of problems does help in many ways, one of which is general problem-solving, even outside of math, acquiring a certain amount of discipline and just feeling good. You're a nice person and I hope you do well in life.

OpenStudy (anonymous):

thanks and u too